OpenStudy (anonymous):

help i need the slope of -5x+2y=4 and -15+6y=0

OpenStudy (bradely):

-5x+2y=4 Add 5x both sides 5x-5x+2y=5x+4 2y=5x+4 Divide each side by y=(5/2)x+2 slope 5/2 -15+6y=0 Add 15 both sides 6y=15 y=15/6 Slope 0 Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

thanks